OpenStudy (anonymous):

Describe how the transformed function below is obtained from its parent function. y=(x-10)^3 a.The parent function y=x^3 is shifted 10 units to the right b.The parent function y=x^3 is shifted 10 units to the left. c.The parent function y=(-x)^3 is strectched vertically by a factor of 10. d.The parent function y=x^3 is shifted 10 units down.

OpenStudy (amistre64):

you can prolly narrow it down to 2 options right away

OpenStudy (amistre64):

nothing is being added to y to move it up or down nothing is multiplying y to stretch it the only thing that is modified is the x values .. which determine right and left positions

OpenStudy (amistre64):

we love parent functions: since they exist at the origin, 0,0 y = x^3: 0 = 0^3 for what value of x, in the modified version, do we acheive this 0,0 format?

OpenStudy (anonymous):

ok, so what would that make it?

OpenStudy (amistre64):

again .... for what value of x, in the modified version, do we acheive this 0,0 format?

OpenStudy (anonymous):

0,0 is the orign

OpenStudy (amistre64):

yes y = x^3 has a nice 0,0 format the modified version y = (x-10)^3 has been moved by x units left or right; so y=0 hasnt changed. just (x-10)^3 = 0 is altered. when does (x-10)^3 = 0 ?

OpenStudy (amistre64):